OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

\(\Large 7sin^2x-14sinx+2=-5\\\Large 7sin^2x-14sinx+7=0\\\Large sin^2x-2sinx+1=0\) Let \(\Large t=sinx\) the equation become: \(\Large t^2-2t+1=0\\\Large \Rightarrow t=1\) Thus, \(\Large sinx=1 \Rightarrow x=\frac{\pi}{2}+2k\pi\ \ (k\in Z) \)